if the expression [sin(x/2) + cos(x/2) - itan(x/2)]/{1+i2sin(x/2)} is real, then the set of all possible values of x is _________

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear raman

[sin(x/2) + cos(x/2) - itan(x/2)]/{1+i2sin(x/2)}

multiple 1-21sin(x/2)  in nominator and denomenator and equate coefficient of i to zero

-tanx/2 -2sinx/2(sinx/2 +cosx/2) =0

or sinx/2 ( 1+ 2cosx/2((sinx/2 +cosx/2)) =0

 or sinx/2(2+sinx/2 +cosx/2)=0

 term in the bracket can not equal to zero

so sinx/2 =0

 x/2 =n∏

x=2n∏  for n any intiger


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin